An American company export product to Japan which sells for 5000 yen. All their costs are incurred in the USA and measured in dollars. The current exchange rate is 102 and their profit margin is 68%. If the exchange rate moves from 102 to 98 but they keep yen price the same, what would be the effect on their profit margin?
Answer #1

Sales=5000 yen=5000/102=49.01960784 dollars

Costs=(1-68%)*49.01960784=15.68627451 dollars

New Sales=5000 yen=5000/98=51.02040816 dollars

New costs=15.68627451 dollars

New Profit Margin=(51.02040816-15.68627451)/51.02040816=69.255%

Profit margin increases by 1.255%